OpenStudy (anonymous):

efflux velocity is proportional to the square root of the distance from the surface of the pol down to the hole. If efflux velocity decreases by a factor of 2, then this means that the distance decreases by a factor of sqrt{2}. Correct or incorrect?

OpenStudy (kainui):

It becomes simpler if we write this out as an equation and plug it in. For instance if you have a cube and decrease the side length by a factor of 2 then how does the volume change? \[V_1=x^3\] just take this equation in my example, and divide our length by 2 \[V_2=(\frac{x}{2})^3 =\frac{x^3}{8}\] but notice x^3 is our original volume! \[V_2=\frac{V_1}{8}\] So that means by decreasing the length of a side of a cube by a half decreases the volume by an eighth! You should be able to apply a similar reasoning here.

OpenStudy (anonymous):

Thank you!